# DCT-CryptoNets: Scaling Private Inference in the Frequency Domain

### v1

@@ Abstract
The convergence of fully homomorphic encryption (FHE) and machine learning offers unprecedented opportunities for private inference of sensitive data. FHE enables computation directly on encrypted data, safeguarding the entire machine learning pipeline, including data and model confidentiality. However, existing FHE-based implementations for deep neural networks face significant challenges in computational cost, latency, and scalability, limiting their practical deployment. This paper introduces DCT-CryptoNets, a novel approach that leverages frequency-domain learning to tackle these issues. Our method operates directly in the frequency domain, utilizing the discrete cosine transform (DCT) commonly employed in JPEG compression. This approach is inherently compatible with remote computing services, where images are usually transmitted and stored in compressed formats. DCT-CryptoNets reduces the computational burden of homomorphic operations by focusing on perceptually relevant low-frequency components. This is demonstrated by substantial latency reduction of up to 5.3$\times$ compared to prior work on image classification tasks, including a novel demonstration of ImageNet inference within 2.5 hours, down from 12.5 hours compared to prior work on equivalent compute resources. Moreover, DCT-CryptoNets improves the reliability of encrypted accuracy by reducing variability (e.g., from ±2.5% to ±1.0% on ImageNet). This study demonstrates a promising avenue for achieving efficient and practical privacy-preserving deep learning on high resolution images seen in real-world applications.
@@ 1 Introduction
Escalating concerns related to privacy and security in machine learning have fueled exploration of private inference techniques. These approaches aim to protect both sensitive data and model parameters while maintaining a high quality user experience, especially when related to inference latency and accuracy. To address these challenges, private inference solutions based on cryptographic principles, such as fully homomorphic encryption (FHE) have been proposed. However, such strong security guarantees often comes with significant computational overhead and latency, creating barriers to widespread adoption.
@@
Early work on fully homomorphic encrypted neural networks (FHENNs) faced limitations in both latency and accuracy. Due to the limited native operations of homomorphic encryption, primarily addition and multiplication, many prior methods resorted to approximating non-linear activation functions using polynomials [8, 4, 7]. However, this approach introduces accuracy degradation as networks deepen, due to the cumulative effect of approximation errors. Newer homomorphic encryption schemes like TFHE (FHE over the Torus) can handle non-linear activation functions without the need for approximations. Even with such features, optimizing the efficiency of homomorphic operations (HOPs), which include convolutions and non-linear activations in the context of neural networks, remains a key area of research.
@@
To reduce the computational burden of HOPs in FHENNs, previous work has investigated strategies for optimizing convolutions in the encrypted domain [22, 19, 16, 29, 31] and designing architectures that minimize the use of non-linear activations [11, 14]. In this work, we prioritize optimizing non-linear activations by operating in the frequency domain. This approach is driven by the significant computational cost of non-linear activations in FHE (e.g., ReLU operations consuming $\sim$32.6of total inference time [19]), and their heightened sensitivity to spatial dimensions. As demonstrated in [14], these activations are disproportionately concentrated in the early layers of networks where spatial dimensionality is the largest (e.g., 48% within the first quarter of a ResNet-18 network).
@@
Figure 1: Scalability of state of the art image classification methods in FHENN. DCT-CryptoNets is able to reduce latency by 5.3$\times$ compared to only other method able to infer on ImageNet (SHE [21]). To ensure a fair comparison, SHE latency values were normalized to the same computational resources as DCT-CryptoNets. CKKS-based methods have difficulty scaling to larger networks and datasets due to their highly approximate nature. [^2] Lee et al. (2) [18] to scale to a ResNet-110 (1.7M parameters). Kim et al [16] scale to a Plain-18 network (ResNet-18 without skip connections) but only encrypt the last 8 layers when running on ImageNet.
@@
While notable advancements have been made, many FHENN optimizations remain tailored to small networks and images (e.g., 32$\times$32). Scaling FHENN to ImageNet and larger networks presents a formidable challenge. Increasing both network size and image dimensions introduces latency from not only convolutions but also non-linear activations, which are "free" in unencrypted networks, and an expensive homomorphic bootstrapping operation [10]. This homomorphic bootstrap is used to "reset" noise in the ciphertext that accumulate for every encrypted operation (addition and multiplication), which can eventually corrupt the ciphertext making decryption impossible. Approximately 31.6% of total inference time is spent on this bootstrapping operation [19]. While attempts have been made to implement FHENN on ImageNet, these efforts highlight the substantial trade-offs involved: either prohibitive latency (e.g., 2.5 days for ResNet-18 [21]) or severely limited encryption scope (e.g., encrypting only 8 layers in ResNet-18 [16]).
@@
This work introduces DCT-CryptoNets, a novel framework that addresses the computational challenges of fully homomorphic encrypted neural networks (FHENNs). We utilize the Discrete Cosine Transform (DCT) to convert image tensors into frequency tensors with 8$\times$8 filters, significantly reducing spatial dimensionality. This not only aligns with the human visual system’s differential sensitivity to perceptually relevant low-frequency information [23] but also mitigates the computational burden of non-linear activations by 34.8%. This yields two additional advantages: (1) it reduces computationally expensive homomorphic bootstrapping, as fewer ReLU activations result in less ciphertext noise accumulation, and (2) it improves the stability of encrypted accuracy through the reduction of bootstrap operations which accumulate approximation errors. Additionally, the emphasis on low-frequency information boosts accuracy compared to traditional RGB-based CNNs due to focusing on visually salient information. Overall, our method demonstrates substantial latency improvements in convolution-based FHENNs across a range of image classification datasets, culminating in a novel demonstration of ImageNet inference within 2.5 hours – a significant advancement over prior work that required an estimated 12.5 hours on equivalent compute resources (previously reported as 2.5 days) [21]. These results underscore the potential of DCT-CryptoNets for accelerating privacy-preserving deep learning applications.
@@
  • We propose DCT-CryptoNets to achieve significant latency improvements in FHENNs. Our focus on low-frequency components reduces the impact of computationally expensive non-linear activations and homomorphic bootstrapping, resulting in a 5.3$\times$ latency reduction on ImageNet, while maintaining or even boosting accuracy compared to RGB-based networks due to the focus on perceptually salient information.
  • We show that DCT-CryptoNets also bolsters the reliability of encrypted accuracy. By mitigating the need for homomorphic bootstrap operations, these optimizations curb error accumulation. This results in a notable reduction in encrypted accuracy variability (e.g., from ±2.5% to ±1.0% on ImageNet).
  • We show that DCT-CryptoNets exhibits superior scalability with increasing resolution. This is evident in the amplified reduction of homomorphic operations (HOPs) observed for larger images (e.g., a 33.1% vs. 44.1% reduction in HOPs when applying DCT-based frequency optimizations to a 224$\times$224 vs. 448$\times$448 image respectively).
@@ 2.1 Crytographic Protocols for Private Inference
Homomorphic encryption relies on lattice-based cryptography [30] and the Learning with Errors (LWE) problem [24], leveraging the computational hardness of solving equations on high-dimensional lattices with added noise. Lattice-based cryptography has gained significant recognition, with 3 out of the 4 finalists in the National Institute of Standards and Technology (NIST) post-quantum cryptography standardization process being based on this approach.
Many state-of-the-art private inference techniques leverage hybrid, "interactive" approaches that combine homomorphic encryption (HE) with other cryptographic protocols such as multi-party computation (MPC) [15, 26, 17]. In MPC, multiple parties jointly compute a function on their private inputs without disclosing those inputs to one another. While ensuring no party learns more than the final result, MPC still can leak information as communication between parties is necessary. Although robust, these methods still pose a security risk in low-trust environments.
In contrast, "non-interactive" FHE-only methods offer superior data confidentiality. FHE ensures that neither raw data nor intermediate values are exposed in plaintext. This makes FHE particularly well-suited for scenarios with highly sensitive data and model parameters, especially in situations where there is limited trust between parties. However, this enhanced privacy comes at the cost of increased computational overhead compared to MPC and hybrid approaches.
@@ 2.2 Limitations of Existing FHENN Schemes
FHE schemes serve as the fundamental cryptographic building blocks for various application-specific methodologies. The choice of an FHE scheme involves a careful evaluation of the trade-offs inherent to each scheme to ensure optimal performance. FHENNs based on earlier schemes such as BFV [9] and BGV [3] were known for their efficiency for encrypted addition and multiplication. However, these methods often faced scalability challenges due to their reliance on polynomial approximations of activations (PAAs) and pooling operations. PAAs introduce accuracy degradation due to the cumulative effect of approximation errors in deeper networks. To mitigate accuracy degradation, techniques to increase the polynomial degree of PAAs have been introduced [13]. Despite this, the computational cost of PAA-based non-linear activations increases exponentially with network depth, becoming a major bottleneck to latency in deeper networks [21]. A more recent scheme, CKKS [5], offers a compelling advantage for neural network applications by operating directly on floating-point numbers. Nevertheless, CKKS inherits the same challenges associated with PAAs and also introduces additional accuracy concerns due to its reliance on approximate arithmetic, which can lead to significant error accumulation in deep networks.
Further efforts to optimize PAA-based methods have shown promise but remain limited. Lee et al. [19] introduced a binary-tree implementation for ReLU in the CKKS scheme, yet scalability is confined to smaller models like ResNet-20. Falcon [22] introduced fast homomorphic discrete Fourier transform (HDFT) for convolutions and fully connected layers in BFV, significantly reducing operations. However, the approach still suffers from limited accuracy (76.5% on CIFAR-10). Kim et al. [16] extended HDFT to CKKS, but only the final 8 layers of a Plain-18 network are encrypted when operating on ImageNet. Lee et al. [18] proposed removing the imaginary components of PAAs, thus enabling scalability up to ResNet-110 (1.7M parameters). While PAA-based methods have shown promise for smaller networks, their inherent constraints in handling increased depth and computational complexity hinder their scalability to deeper networks.
@@ 2.3 Advantages of TFHE
In contrast to other HE schemes, TFHE (FHE over the Torus) operates within the domain of the torus modulo 1, representing ciphertexts as elements of this structure. TFHE excels in performing fast and exact binary operations on encrypted bits, utilizing integer and logic gates. The foundation for these operations lies in the M-th cyclotomic polynomial, denoted as $\Phi(X)$. Its degree, represented by $N$, is typically chosen as a power of 2 for efficiency reasons, and simplifies $\Phi(X)$ to $X^{N}+1$. By defining polynomial rings over real coefficients $R_{N}[X]:=\mathbb{R}[X]/(X^{N}+1)$ and integer coefficients $Z_{N}[X]:=\mathbb{Z}[X]/(X^{N}+1)$, we establish the Torus polynomial ring $\mathbb{T}_{N}[X]:=R_{N}[X]/Z_{N}[X]=\mathbb{T}[X]/(X^{N}+1)$ which forms a $\mathbb{Z}_{N}[X]$-module. This algebraic structure is fundamental to TFHE, facilitating both addition and external multiplication by polynomials of $\mathbb{Z}_{N}[X]$.
Unlike HE schemes that rely on PAAs, TFHE’s ability to directly implement non-linear activations using Boolean/integer arithmetic is a key factor in scalability. Directly implementing non-linear activations avoids the accuracy degradation often associated with PAAs in deeper networks. SHE (Lou and Jiang [21]), a TFHE-based FHENN, is able to show scalability to ImageNet. They do so by employing a bit-series representation and techniques like logarithmic weight quantization and bit-shift-based convolutions. However, SHE still suffers from prohibitive latency on ImageNet (e.g., ResNet-18 in 2.5 days, ShuffleNet in 5 hours per image). Furthermore, SHE utilizes a leveled TFHE scheme which necessitates higher multiplicative depths and larger ciphertexts for deeper networks.
In our approach, we apply programmable bootstrapping mechanisms (PBS) [6] to support arbitrarily deep neural networks. PBS serves two crucial purposes: (1) better ciphertext noise reduction by reordering rotation and key-switch ciphertext operations and (2) enabling homomorphic evaluation of any function expressible as a lookup-table. PBS’s ability to homomorphically evaluate functions expressed as lookup-tables makes it well-suited for implementing non-linear activation functions.

### v2

@@ Abstract
The convergence of fully homomorphic encryption (FHE) and machine learning offers unprecedented opportunities for private inference of sensitive data. FHE enables computation directly on encrypted data, safeguarding the entire machine learning pipeline, including data and model confidentiality. However, existing FHE-based implementations for deep neural networks face significant challenges in computational cost, latency, and scalability, limiting their practical deployment. This paper introduces DCT-CryptoNets, a novel approach that operates directly in the frequency-domain to reduce the burden of computationally expensive non-linear activations and homomorphic bootstrap operations during private inference. It does so by utilizing the discrete cosine transform (DCT), commonly employed in JPEG encoding, which has inherent compatibility with remote computing services where images are generally stored and transmitted in this encoded format. DCT-CryptoNets demonstrates a substantial latency reductions of up to 5.3$\times$ compared to prior work on benchmark image classification tasks. Notably, it demonstrates inference on the ImageNet dataset within 2.5 hours (down from 12.5 hours on equivalent 96-thread compute resources). Furthermore, by learning perceptually salient low-frequency information DCT-CryptoNets improves the reliability of encrypted predictions compared to RGB-based networks by reducing error accumulating homomorphic bootstrap operations. DCT-CryptoNets also demonstrates superior scalability to RGB-based networks by further reducing computational cost as image size increases. This study demonstrates a promising avenue for achieving efficient and practical private inference of deep learning models on high resolution images seen in real-world applications. Our code can be found at https://github.com/ar-roy/dct-cryptonets.
@@ 1 Introduction
Escalating privacy and security concerns in machine learning have fueled exploration of private inference techniques. These approaches aim to protect both sensitive data and model confidentiality while maintaining a high quality user experience, especially when related to inference latency and accuracy. To address these challenges, private inference solutions based on cryptographic principles, such as fully homomorphic encryption (FHE) have been proposed. However, FHE’s strong security guarantee often comes with significant computational overhead and latency, creating barriers to widespread adoption.
@@
Early work on fully homomorphic encrypted neural networks (FHENNs) faced limitations in both latency and accuracy. Due to the limited native operations of homomorphic encryption, primarily addition and multiplication, many prior methods resorted to approximating non-linear activation functions using polynomials (Dowlin et al., 2016; Brutzkus et al., 2018; Chou et al., 2018). However, this approach introduces accuracy degradation as networks deepen, due to the cumulative effect of approximation errors. Newer homomorphic encryption schemes like TFHE (FHE over the Torus) (Chillotti et al., 2019) can handle non-linear activation functions without the need for approximations. Still, optimizing the efficiency of homomorphic operations (HOPs), which include convolutions and non-linear activations in the context of neural networks, remains a key area of research.
@@
To reduce the computational burden of HOPs in FHENNs, previous work has investigated strategies for optimizing convolutions in the encrypted domain (Lou et al., 2020; Lee et al., 2022b; Kim & Guyot, 2023; Ran et al., 2023; Rovida & Leporati, 2024) and designing architectures that minimize the use of non-linear activations (Ghodsi et al., 2020; Jha et al., 2021). In this work, we prioritize optimizing non-linear activations by operating in the frequency domain. This approach is motivated by the significant computational cost of non-linear activations in FHE (e.g., ReLU operations consuming $\sim$32.6of total inference time (Lee et al., 2022b)), and the fact that non-linear activations scale primarily with the spatial extent of an input. As demonstrated in Jha et al. (2021), non-linear activations are disproportionately concentrated in the early layers of networks where spatial dimensionality is the largest (e.g., 48% within the first quarter of a ResNet-18 network).
@@
Figure 1: Scalability of state of the art image classification methods in FHENN. DCT-CryptoNets is able to reduce latency by up to 5.3$\times$ compared to SHE (Lou & Jiang, 2019), the only other published method able to infer on ImageNet. To ensure a fair comparison, SHE latency values were normalized to the same computational resources as DCT-CryptoNets (96-threads). CKKS-based methods have difficulty scaling to larger networks and datasets due to their highly approximate nature. [^2] Lee et al. (2022a) scale to ResNet-110 (1.7M parameters). Kim & Guyot (2023) scale to a Plain-18 network (ResNet-18 without skip connections) but only encrypt the last 8 layers when running on ImageNet.
@@
While notable advancements have been made, many FHENN methods remain tailored to small networks and images (e.g., 32$\times$32). Scaling FHENN to ImageNet and larger networks presents a formidable challenge. Increasing both network size and image resolution introduces latency from not only convolutions but also non-linear activations, which are "free" in unencrypted networks, and an expensive homomorphic bootstrapping operation (Gentry, 2010). Homomorphic bootstrap is used to "reset" noise in the ciphertext that accumulate for every encrypted operation (addition and multiplication), which can eventually corrupt the ciphertext making decryption impossible. Approximately 31.6% of total inference time is spent on such bootstrapping operations (Lee et al., 2022b). While attempts have been made to implement FHENN on large-scale datasets such as ImageNet, these efforts highlight the substantial trade-offs involved: either prohibitive inference latency (e.g., 2.5 days for ResNet-18 in Lou & Jiang (2019)) or severely limited encryption scope (e.g., encrypting only 8 layers in ResNet-18 in Kim & Guyot (2023)) as shown in Footnote 2.
@@
This work introduces DCT-CryptoNets, a novel framework that addresses the computational challenges of fully homomorphic encrypted neural networks (FHENNs). Traditional convolutional neural networks operate on raw pixel data, learning features from spatial intensity variations. Instead, we utilize Discrete Cosine Transforms (DCT) to represent images in the frequency domain, enabling our models to learn features from the rate of change in intensities (Gueguen et al., 2018; Ehrlich & Davis, 2019). This not only aligns with the human visual system’s differential sensitivity to perceptually relevant low-frequency information (Luo et al., 2022), but the reduction in spatial dimensionality by operating in the frequency domain also mitigates the computational burden of non-linear activations. DCT-CryptoNets yields two additional advantages: (1) it reduces computationally expensive homomorphic bootstrapping, as fewer non-linear activations result in less ciphertext noise accumulation, and (2) it improves the reliability of encrypted accuracy through the reduction of bootstrap operations which accumulate approximation errors. Additionally, the emphasis on low-frequency information has the potential to boost accuracy compared to traditional RGB-based CNNs due to focusing on visually salient information. Overall, our method demonstrates substantial latency improvements in convolution-based FHENNs across a range of image classification datasets, culminating in a novel demonstration of ImageNet inference within 2.5 hours – a significant advancement over SHE (Lou & Jiang, 2019) that required 12.5 hours on equivalent compute resources (previously reported as 2.5 days). These results underscore the potential of DCT-CryptoNets for accelerating privacy-preserving deep learning applications.
@@
  • We propose DCT-CryptoNets to achieve significant latency improvements in image classification FHENNs by utilizing DCT. Our focus on low-frequency components reduces the impact of computationally expensive non-linear activations and homomorphic bootstrapping, resulting in a 5.3$\times$ latency reduction on ImageNet, while maintaining or even boosting accuracy compared to RGB-based networks due to the focus on perceptually salient information.
  • We show that DCT-CryptoNets also bolsters the reliability of encrypted accuracy. By curtailing the need for homomorphic bootstrap operations, these optimizations curb error accumulation. This results in a notable reduction in encrypted accuracy variability (e.g., from ±2.5% to ±1.0% on ImageNet).
  • We show that DCT-CryptoNets exhibits superior scalability with increasing image resolution. This is evident in the amplified reduction of homomorphic operations (HOPs) observed for larger images (e.g., a 33.1% vs. 44.1% reduction in HOPs when applying DCT-based frequency optimizations to a 224$\times$224 vs. 448$\times$448 image respectively).
@@ 2.1 Crytographic Protocols for Private Inference
Homomorphic encryption relies on lattice-based cryptography (Regev, 2005) and the Learning with Errors (LWE) problem (Lyubashevsky et al., 2010), leveraging the computational hardness of solving equations on high-dimensional lattices with added noise. Lattice-based cryptography has gained significant traction, with 3 out of the 4 finalists in the National Institute of Standards and Technology (NIST) post-quantum cryptography standardization being based on this approach.
Many state-of-the-art private inference techniques leverage hybrid, "interactive" approaches that combine homomorphic encryption with other cryptographic protocols such as multi-party computation (MPC) (Juvekar et al., 2018; Mishra et al., 2020; Knott et al., 2021). In MPC, multiple parties jointly compute a function on their private inputs without disclosing those inputs. While ensuring no party learns more than the final result, MPC still can leak information as communication between parties is necessary. Although robust, these methods still pose a security risk in low-trust environments. In contrast, "non-interactive" FHE-only methods offer superior data confidentiality. FHE ensures that neither raw data nor intermediate values are exposed in plaintext. This makes FHE particularly well-suited for scenarios with highly sensitive data and model parameters, especially in situations where there is severely limited trust between parties (see Section A.1). However, this enhanced privacy comes at the cost of increased computational overhead compared to MPC and hybrid approaches.
@@ 2.2 Limitations of Existing FHENN Schemes
FHE schemes serve as the fundamental cryptographic building blocks for various application-specific methodologies. The choice of an FHE scheme involves a careful evaluation of the trade-offs inherent to each scheme to ensure optimal performance. FHENNs based on earlier schemes such as BFV (Fan & Vercauteren, 2012) and BGV (Brakerski et al., 2012) were known for their efficiency for encrypted addition and multiplication. However, these methods often faced scalability challenges due to their reliance on polynomial approximations of activations (PAAs) and pooling operations. This reliance on PAAs introduces accuracy degradation due to the cumulative effect of approximation errors in deeper networks. To mitigate accuracy degradation, techniques to increase the polynomial degree of PAAs have been introduced (Hesamifard et al., 2019). Despite this, the computational cost of PAA-based non-linear activations increases exponentially with network depth, becoming a major bottleneck to latency in deeper networks (Lou & Jiang, 2019). A more recent scheme, CKKS (Cheon et al., 2017), offers a compelling advantage for neural network applications by operating directly on floating-point numbers and supporting single instruction multiple data (SIMD) operations. Nevertheless, CKKS inherits the same challenges associated with PAAs and also introduces additional accuracy concerns due to its reliance on approximate arithmetic, which can lead to significant error accumulation in deep networks.
Further efforts to optimize PAA-based methods have shown promise but remain limited. Lee et al. (2022b) introduced a binary-tree implementation for ReLU in the CKKS scheme, yet scalability is confined to smaller models like ResNet-20. Falcon (Lou et al., 2020) introduced fast homomorphic discrete Fourier transform (HDFT) for convolutions and fully connected layers in BFV, significantly reducing operations. However, this approach still suffers from limited accuracy (76.5% on CIFAR-10) and requires specialized convolution blocks, hindering broader applicability with existing CNN architectures. Kim & Guyot (2023) extended HDFT to CKKS, but only the final 8 layers of a Plain-18 network are encrypted when operating on ImageNet. Lee et al. (2022a) proposed removing the imaginary components of PAAs, thus enabling scalability up to ResNet-110 (1.7M parameters). While PAA-based methods have shown promise for smaller networks, their inherent constraints in handling increased depth and computational complexity hinder their scalability to deeper networks.
@@ 2.3 Advantages of TFHE
In contrast to other HE schemes, TFHE (FHE over the Torus) (Chillotti et al., 2019) operates within the domain of the torus modulo 1, representing ciphertexts as elements of this structure. TFHE excels in performing fast and exact binary operations on encrypted bits, utilizing integer and logic gates. The foundation for these operations lies in the M-th cyclotomic polynomial, denoted as $\Phi(X)$. Its degree, represented by $N$, is typically chosen as a power of 2 for efficiency reasons, and simplifies $\Phi(X)$ to $X^{N}+1$. By defining polynomial rings over real coefficients $R_{N}[X]:=\mathbb{R}[X]/(X^{N}+1)$ and integer coefficients $Z_{N}[X]:=\mathbb{Z}[X]/(X^{N}+1)$, we establish the Torus polynomial ring $\mathbb{T}_{N}[X]:=R_{N}[X]/Z_{N}[X]=\mathbb{T}[X]/(X^{N}+1)$ which forms a $\mathbb{Z}_{N}[X]$-module. This algebraic structure is fundamental to TFHE, facilitating both addition and external multiplication by polynomials of $\mathbb{Z}_{N}[X]$.
Unlike HE schemes that rely on PAAs, TFHE’s ability to directly implement non-linear activations using Boolean/integer arithmetic is a key factor in scalability. Directly implementing non-linear activations avoids the accuracy degradation often associated with PAAs in deeper networks. SHE (Lou & Jiang, 2019), a TFHE-based FHENN, shows scalability to ImageNet. They do so by employing a bit-series representation and techniques like logarithmic weight quantization and bit-shift-based convolutions. However, SHE still suffers from prohibitive latency on ImageNet (e.g., ResNet-18 in 2.5 days, ShuffleNet in 5 hours per image). Furthermore, SHE utilizes a leveled TFHE scheme which necessitates higher multiplicative depths to avoid decryption failures as a network deepens. This higher multiplicative depth budget results in larger ciphertexts which consequently increases latency in these deeper networks.
In our approach, we apply programmable bootstrapping (PBS) mechanisms (Chillotti et al., 2021) to support arbitrarily deep neural networks and mitigate the limitations of leveled schemes. PBS serves two crucial purposes: (1) better ciphertext noise reduction by reordering rotation and key-switch ciphertext operations and (2) enabling homomorphic evaluation of any function expressible as a lookup-table. PBS’s ability to homomorphically evaluate functions expressed as lookup-tables makes it well-suited for implementing non-linear activations.